OpenStudy (anonymous):

Choose the correct simplification of the expression a12 b8/ a2 a4

OpenStudy (anonymous):

A a10b4 B a6b2

OpenStudy (anonymous):

Choose the correct simplification of the expression (see attached photo for expression)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (anonymous):

when you are dividing expressions with exponents you subtract the exponents. And inversely when you multiply expressions with exponents you add them

OpenStudy (anonymous):

B is not correct what is your second choice

OpenStudy (anonymous):

I'm sorry A is the correct choice

OpenStudy (anonymous):

12-2 is 10 and 8-4 is 4

OpenStudy (anonymous):

@MATTW20 thank you so much could you help me with another?

OpenStudy (anonymous):

most definitely

OpenStudy (anonymous):

ready when you are

OpenStudy (anonymous):

Choose the correct simplification of the expression (see attached photo for expression)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

should be \[\frac{ a ^{6} }{ b ^{14} }\]

OpenStudy (anonymous):

so its going to be C

OpenStudy (anonymous):

yes when expressions with exponents are raise to a power you multiply the powers so it would be \[\frac{ a ^{3*2} }{ b ^{7*2}}\] which gets you \[\frac{ a ^{6} }{ b ^{14} }\]

OpenStudy (anonymous):

@MATTW20 could you help me with one more?

OpenStudy (anonymous):

Choose the correct simplification of the expression (see attached photo)

OpenStudy (anonymous):

Are you understanding this? I can help you but I want to help you so that you can do it on your own eventually.

OpenStudy (anonymous):

@MATTW20 yes i'm understanding it

OpenStudy (anonymous):

A

OpenStudy (anonymous):

\[4ba ^{10}\]

OpenStudy (anonymous):

B 4a10b D

OpenStudy (anonymous):

B

OpenStudy (anonymous):

okay great thanks so much for all your help @MATTW20 :) YOU ARE A LIFESAVER

OpenStudy (anonymous):

\[4a ^{10}b\]